Problem 7-71B (Algorithmic)
Depreciation Schedules

Dunn Corporation acquired a new depreciable asset for $142,000. The asset has a 5-year expected life and a residual value of zero.

Required:

1. Prepare a depreciation schedule for all 5 years of the asset's expected life using the straight-line depreciation method. If an amount is zero, enter "0".

Dunn Corporation
Straight-Line Depreciation Method
Five Years
End of Year Depreciation Expense Accumulated Depreciation Book Value
$
Year 1 $ $
Year 2
Year 3
Year 4
Year 5

2. Prepare a depreciation schedule for all 5 years of the asset's expected life using the double-declining-balance depreciation method. Round all answers to the nearest dollar. If an amount is zero, enter "0".

Dunn Corporation
Double-Declining-Balance Depreciation Method
Five Years
End of Year Depreciation Expense Accumulated Depreciation Book Value
Year 1
Year 2
Year 3
Year 4
Year 5
Answer #1

Solution

Dunn Corporation
Straight-Line Depreciation Method
Five Years
End of Year Depreciation Expense Accumulated Depreciation Book Value
$     142,000
Year 1 $       28,400 $           28,400 $     113,600
Year 2 $       28,400 $           56,800 $       85,200
Year 3 $       28,400 $           85,200 $       56,800
Year 4 $       28,400 $         113,600 $       28,400
Year 5 $       28,400 $         142,000 $              -  

.

Dunn Corporation
Double-Declining-Balance Depreciation Method
Five Years
End of Year Depreciation Expense Accumulated Depreciation Book Value
$     142,000
Year 1 $       56,800 $           56,800 $       85,200
Year 2 $       34,080 $           90,880 $       51,120
Year 3 $       20,448 $         111,328 $       30,672
Year 4 $       12,269 $         123,597 $       18,403
Year 5 $       18,403 $         142,000 $              -  

Working

Straight line Method
A Cost $ 142,000
B Residual Value $ 0
C=A - B Depreciable base $ 142,000
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $ 28,400

.

Double declining Method
A Cost $ 142,000
B Residual Value $ 0
C=A - B Depreciable base $ 142,000
D Life [in years] 5
E=C/D Annual SLM depreciation $ 28,400
F=E/C SLM Rate 20.00%
G=F x 2 DDB Rate 40.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $ 142,000 40.00% $ 56,800 $ 56,800 $ 85,200
2 $ 85,200 40.00% $ 34,080 $ 90,880 $ 51,120
3 $ 51,120 40.00% $ 20,448 $ 111,328 $ 30,672
4 $ 30,672 40.00% $ 12,269 $ 123,597 $ 18,403
5 $ 18,403 40.00% $ 18,403* $ 142,000 $ 0

*In year 5 depreciation is not 40% instead it is adjusted to make ending book value equal to zero.